Given g(n)-2n-5;f(n)=3n-4. Find (g°f)(-2n)


1
Expert's answer
2021-02-03T04:52:03-0500

Given g(n)-2n-5;f(n)=3n-4. Find (g°f)(-2n)


Solution:

"(g \circ f)(-2n)=g(f(-2n))=g(3\cdot(-2n)-4)=g(-6n-4)="

"-2(-6n-4)-5=12n+3"

Answer: "(g\circ f)(-2n)=12n+3" .